# A Cold and Super-Puffy Planet on a Polar Orbit

### v1

@@ 6.5 Explaining the inferred low density
Although sub-day rotation periods may be reasonable, or even expected, for a Neptune-like planet, the required values for rocky planets may approach break-up rotation period (e.g., $P_{\rm rot}\lesssim 1$ h for $R_{p}\lesssim 1.1\,R_{\oplus}$). Assuming instead a Neptune-like planet, the rotation must remain rapid despite tidal damping. The corresponding spin-down timescale is (e.g., A. C. M. Correia & J. Laskar 2010)
@@
In summary, producing a deep transit with circumplanetary rings is primarily constrained by Poynting-Robertson drag, which can make rings unsustainable at the observed temperatures. In comparison, the rotation periods and tidal dissipation needed to sustain a high obliquity are less restrictive.
@@ 7 Conclusions
We have reported the discovery and characterization of TOI-4507 b, a transiting cold Neptune on a nearly polar orbit. With a radius of $8.22\pm 0.08\,R_{\oplus}$, a mass $30\,M_{\oplus}$, and a bulk density <0.3<0.3 g/cm3, TOI-4507 b is among the longest-period ($P\sim 105$ d) super-puffs discovered to date. Its nearly polar orbit with a 3D stellar obliquity of $\psi=82.0_{-2.4}^{+2.6}$ deg makes it one of the youngest polar planets known and the longest-period system with a measured 3D obliquity.
@@
The extreme properties of TOI-4507 b raise several important implications. First, its unusually low density highlights the need to consider both compositional and structural explanations for super-puff planets, including massive H/He envelopes, inflation mechanisms, or obscuration by circumplanetary material. However, our analysis suggests that long-lived opaque rings are unlikely to survive at the planet’s equilibrium temperature. Second, the nearly polar and long-period orbit suggests that the orbital plane was either primordially tilted or secularly excited by an outer companion. The detection of a long-term RV trend further hints at additional bodies that may have influenced the system’s architecture.
@@
Given its brightness, young age, and extremely low density, TOI-4507 b is a prime target for follow-up studies. Transmission spectroscopy with JWST can test the super-puff hypothesis by probing its atmospheric composition, while continued RV and astrometric monitoring can confirm the presence of an outer companion and further constrain the dynamical history of the system. TOI-4507 b thus joins a small but growing population of polar Neptunes and super-puffs, providing a valuable laboratory for understanding the formation, evolution, and diversity of planetary systems.
@@
We would like to thank Yubo Su for suggesting the rotational constraint on the sizes of the rings. We also would like to thank Songhu Wang, Xian-Yu Wang, and the Red Worlds Lab group at the University of Amsterdam for their useful discussions and suggestions that strengthened this manuscript. J.I.E.-R. gratefully acknowledges support from the John and A-Lan Reynolds Faculty Research Fund, from the ANID BASAL project FB210003, and from the ANID Doctorado Nacional grant 2021-21212378. R.B. acknowledges support from FONDECYT Project 1241963 and from ANID – Millennium Science Initiative – ICN12_009. Support for M.C. is provided by ANID’s Millennium Science Initiative through grants ICN12_12009 and AIM23-001, awarded to the Millennium Institute of Astrophysics (MAS); by ANID/FONDECYT Regular grant 1231637; and by ANID/Basal (CATA) grant FB21003. $C.$ acknowledges support from ANID through FONDECYT post-doctoral grant number 3230283. A.V.F. acknowledges the support of the Institute of Physics through the Bell Burnell Graduate Scholarship Fund. This paper was based on observations collected at the European Southern Observatory under ESO programmes 0104.C-0413(A), 106.21ER.001, 112.25W1.001, 114.27CS.001, and 115.286G.001. This paper also made use of data collected by the TESS mission and are publicly available from the Mikulski Archive for Space Telescopes (MAST) operated by the Space Telescope Science Institute (STScI). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This work makes use of observations from the ASTEP telescope. ASTEP benefited from the support of the French and Italian polar agencies IPEV and PNRA in the framework of the Concordia station program and from OCA, INSU, Idex UCAJEDI (ANR- 15-IDEX-01) and ESA through the Science Faculty of the European Space Research and Technology Centre (ESTEC). This research also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 803193/BEBOP), from the Science and Technology Facilities Council (STFC; grant No. ST/S00193X/1, ST/W002582/1 and ST/Y001710/1), and from the ERC/UKRI Frontier Research Guarantee programme (CandY/ EP/Z000327/1). This work has been carried out within the framework of the National Centre of Competence in Research PlanetS supported by the Swiss National Science Foundation under grants 51NF40_182901 and 51NF40_205606. This paper is also based on observations obtained at the SOAR telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP) and NASA Exoplanet Archive websites, which are operated by the California Institute of Technology under contract with NASA under the Exoplanet Exploration Program.

### v2

@@ 7.3 Explaining the Inferred Low Density
Figure 4: Sky-projected obliquity $\lambda$ as a function of the orbital period (left) and age of the system (right). The population of Neptunes/sub-Saturns ($10<M_{p}/M_{\oplus}<90$) is shown in blue. Super-puff planets (Neptunes/sub-Saturns with $\rho_{p}<0.3$ g/cm3) are shown in pink, with TOI-4507 highlighted as a star with a red edge. Other giant planet systems are shown as faint black points. Data from TEPCat with ages from the NASA Exoplanet Archive.
@@
Although sub-day rotation periods may be reasonable, or even expected, for a Neptune-like planet, the required values for rocky planets may approach break-up rotation period (e.g., $P_{\rm rot}\lesssim 1$ h for $R_{p}\lesssim 1.1\,R_{\oplus}$). Assuming instead a Neptune-like planet, the rotation must remain rapid despite tidal damping. The corresponding spin-down timescale is (e.g., A. C. M. Correia & J. Laskar 2010):
@@ 8 Conclusions
We have reported the discovery and characterization of TOI-4507 b, a transiting cold Neptune on a orbit. With a radius of $8.2\pm 0.1\,R_{\oplus}$, a mass $20\,M_{\oplus}$, and a bulk density <0.2<0.2 g/cm3, TOI-4507 b is among the longest-period ($P\sim 105$ d) super-puffs discovered to date.
@@
The properties of TOI-4507 b raise important implications. Its unusually low density highlights the need to consider both compositional and structural explanations for super-puff planets, including massive H/He envelopes, inflation mechanisms, or obscuration by circumplanetary material. However, our analysis suggests that long-lived opaque rings are unlikely to survive at the planet’s equilibrium temperature.
@@
Given its brightness, young age, and extremely low density, TOI-4507 b is a prime target for follow-up studies. Transmission spectroscopy with JWST can test the super-puff hypothesis by probing its atmospheric composition, while continued RV and astrometric monitoring can confirm the presence of an outer companion and further constrain the dynamical history of the system.
TOI-4507 thus joins a small but growing population of providing a valuable laboratory for understanding the formation, evolution, and diversity of planetary systems.
@@
We would like to thank Yubo Su for suggesting the rotational constraint on the sizes of the rings. We also thank Songhu Wang, Xian-Yu Wang, and the Red Worlds Lab group at the University of Amsterdam for their useful discussions and suggestions that strengthened this manuscript. J.I.E.-R. gratefully acknowledges support from the John and A-Lan Reynolds Faculty Research Fund, from the ANID BASAL project FB210003, and from the ANID Doctorado Nacional grant 2021-21212378. R.B. acknowledges support from FONDECYT Project 1241963 and from ANID – Millennium Science Initiative – ICN12_009. Support for M.C. is provided by ANID’s Millennium Science Initiative through grants ICN12_12009 and AIM23-001, awarded to the Millennium Institute of Astrophysics (MAS); by ANID/FONDECYT Regular grant 1231637; and by ANID/Basal (CATA) grant FB21003. $C.$ acknowledges support from ANID through FONDECYT post-doctoral grant number 3230283. A.V.F. acknowledges the support of the Institute of Physics through the Bell Burnell Graduate Scholarship Fund. This paper was based on observations collected at the European Southern Observatory under ESO programmes 0104.C-0413(A), 106.21ER.001, 112.25W1.001, 114.27CS.001, and 115.286G.001. This paper also made use of data collected by the TESS mission and are publicly available from the Mikulski Archive for Space Telescopes (MAST) operated by the Space Telescope Science Institute (STScI). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This work makes use of observations from the ASTEP telescope. ASTEP benefited from the support of the French and Italian polar agencies IPEV and PNRA in the framework of the Concordia station program and from OCA, INSU, Idex UCAJEDI (ANR- 15-IDEX-01) and ESA through the Science Faculty of the European Space Research and Technology Centre (ESTEC). This research also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 803193/BEBOP), from the Science and Technology Facilities Council (STFC; grant No. ST/S00193X/1, ST/W002582/1 and ST/Y001710/1), and from the ERC/UKRI Frontier Research Guarantee programme (CandY/ EP/Z000327/1). This work has been carried out within the framework of the National Centre of Competence in Research PlanetS supported by the Swiss National Science Foundation under grants 51NF40_182901 and 51NF40_205606. This paper is also based on observations obtained at the SOAR telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP) and NASA Exoplanet Archive websites, which are operated by the California Institute of Technology under contract with NASA under the Exoplanet Exploration Program.
Appendix A Radial Velocity Measurements
[…]
Table 4: Radial velocity measurements from HARPS, FEROS, and CORALIE.
BJD RV (m/s) $\sigma_{\rm RV}$ (m/s) Instrument
2458717.932107 26094.5 11.0 FEROS
2458718.898281 26072.7 9.0 FEROS
2458722.879818 26053.6 9.3 FEROS
 $\cdots$  $\cdots$  $\cdots$  $\cdots$
2460667.658069 26062.9 2.3 HARPS
2460692.667466 26060.0 3.8 HARPS
2460744.564697 26059.6 3.3 HARPS
Note. — Table 4 is published in its entirety in the machine-readable format. A portion is shown here for guidance regarding its form and content.
Appendix B Source of the Photometric Variability
Figure 5: Observed modulation in the light curves of TOI-4507 (upper panels) and the nearby source Gaia DR3 4657949658179626624 (bottom panels). a) Lomb-Scargle periodogram of the 2-min cadence TESS light curve of TOI-4507. A false alarm probability (FAP) level of 1% is marked with a red dashed line. The black dashed line marks the highest-power peak, indicating a period of 1.7 days. b) TESS light curve phased to the period of 1.7 days. All the work done here made use of the 2-minute cadence light curve, but for illustrative purposes only, the 30-minute cadence TESS light curve is shown here in blue, with binned data in red. c) Lomb-Scargle periodogram of the OGLE light curve of Gaia DR3 4657949658179626624, also showing a clear period of 1.7 days. d) OGLE light curve of Gaia DR3 4657949658179626624 phased to the period of 1.7 days.
Appendix C Corner Plot
Figure 6: Joint posterior distributions and histograms of the posteriors for some of the parameters in our RM effect model.